# Tractable and Provably Efficient Distributional Reinforcement Learning with General Value Function Approximation

### v1

@@ Appendix C Proof
Our proof is mainly based on the proof techniques of Rowland et al. [41] and we describe in an extended form. Since their proof also considers the discounted setting, we will define $\mathcal{B}_{r,\gamma}(x)=r+\gamma x$ for discount factor $\gamma\in[0,1)$. By assumption of Bellman closedness, $\psi_{n}\Big{(}(\mathcal{B}_{r,\gamma})_{\#}\bar{\eta}(s^{\prime})\Big{)}$ will be written as $g(r,\gamma,\psi_{1:N}(\bar{\eta}(s^{\prime}))$ for some $g$. By assumption of Bellman unbiasedness and Lemma 4.6, both $\psi_{1:N}(\bar{\eta}(s^{\prime}))$ and $\psi_{n}\Big{(}(\mathcal{B}_{r,\gamma})_{\#}\bar{\eta}(s^{\prime})\Big{)}$ are affine as functions of the distribution $\bar{\eta}(s^{\prime})$,
@@
According to Engert [18], for any translation invariant finite-dimensional space is spanned by a set of function of the form
@@
$\geq\|f-f_{\bar{\eta}^{\prime}}\|^{2}_{\mathcal{Z}_{h}^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-4N(H+1)}-\Big{|}\sum_{n=1}^{N}\sum_{s_{h}^{\tau},a_{h}^{\tau}\in\mathcal{Z}^{k}_{h,\bar{\eta}^{\prime}}}\chi^{\tau}_{h}(f^{(n)})\Big{|}.$
@@
$\|f^{(n)}_{\bar{\eta}}-f^{(n)}_{\bar{\eta}^{\prime}}\|_{\infty}$ $=\max_{s,a}\Big{|}\sum_{n^{\prime}=1}^{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{n^{\prime}}}[\psi_{n^{\prime}}([\mathbb{P}\bar{\eta}](s,a))-\psi_{n^{\prime}}([\mathbb{P}\bar{\eta}^{\prime}](s,a))]r^{n-n^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}/H^{n-1}}\Big{|}$
@@
$=\sum_{n=1}^{N}2(f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau}))(f^{(n)}_{\bar{\eta}}(s_{h}^{\tau},a_{h}^{\tau})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\psi_{n}\Big{(}(\mathcal{B}_{r_{h}^{\tau}})_{\#}\mathbb{E}_{s_{h+1}^{\tau}\sim\mathbb{P}_{h}(\cdot|s_{h}^{\tau},a_{h}^{\tau})}[{\bar{\eta}}(s_{h+1}^{\tau})]\Big{)})}$
@@
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max_{n\in[N]}\Big{\{}2(f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-\psi_{n}\Big{(}(\mathcal{B}_{r_{h}^{\tau}})_{\#}{\bar{\eta}}(s_{h+1}^{\tau})\Big{)})\Big{\}}}\sum_{n=1}^{N}\Big{|}f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})\Big{|}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq 2(H+1)\sum_{n=1}^{N}\Big{|}f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})\Big{|}}$
@@
Lemma C.1.
Let $Q^{k}_{h}(s,a)\coloneqq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\{H,\tilde{f}^{k}_{h}(s,a)+b^{k}_{h}(s,a)\}}$ for some bonus function $b^{k}_{h}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. If $b^{k}_{h}(s,a)\geq w^{(1)}((\mathcal{F}^{N})^{k}_{h},s,a)$ , then with probability at least $21-\delta/$ 2,
@@
Lemma C.2 (Regret decomposition).
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\bm{1}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b^{k}_{h}(s^{k}_{h},a^{k}_{h})}>\epsilon\}\leq\left(\frac{4\beta(\mathcal{F}^{N},\delta)}{\epsilon^{2}}+1\right)\text{dim}_{E}(\mathcal{F}^{N},\epsilon)$
@@
$4\beta(\mathcal{F}^{N},\delta)\geq\|f-g\|^{2}_{\mathcal{Z}^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\geq\|f^{(1)}-g^{(1)}\|^{2}_{\mathcal{Z}^{k}}}\geq L\epsilon^{2}$
@@
Lemma C.4 (Refined version of Lemma 10 in Wang et al. [50]).
@@
We first sort the sequence $\{b^{k}_{h}(s^{k}_{h},a^{k}_{h})\}_{(k,h)\in[K]\times[H]}$ in a decreasing order and denote it by $\{e_{1},\ldots,e_{T}\}(e_{1}\geq e_{2}\geq\dots\geq e_{T})$. By Lemma C.3, for any constant $M>0$ and $e_{t}\geq 1/\sqrt{M}T$, we have
@@
Conditioning on the above event and Lemma C.4, we have

### v2

@@ Appendix D Proof
Our proof is mainly based on the proof techniques of Rowland et al. (2019) and we describe in an extended form. Since their proof also considers the discounted setting, we will define $\mathcal{B}_{r,\gamma}(x)=r+\gamma x$ for discount factor $\gamma\in[0,1)$. By assumption of Bellman closedness, $\psi_{n}\Big{(}(\mathcal{B}_{r,\gamma})_{\#}\bar{\eta}(s^{\prime})\Big{)}$ will be written as $g(r,\gamma,\psi_{1:N}(\bar{\eta}(s^{\prime}))$ for some $g$. By assumption of Bellman unbiasedness and Lemma 3.5, both $\psi_{1:N}(\bar{\eta}(s^{\prime}))$ and $\psi_{n}\Big{(}(\mathcal{B}_{r,\gamma})_{\#}\bar{\eta}(s^{\prime})\Big{)}$ are affine as functions of the distribution $\bar{\eta}(s^{\prime})$,
@@
According to Engert (1970), for any translation invariant finite-dimensional space is spanned by a set of function of the form
@@
$\geq\|f-f_{\bar{\eta}^{\prime}}\|^{2}_{\mathcal{Z}_{h}^{k}}{-4N(H+1)}-\Big{|}\sum_{n=1}^{N}\sum_{s_{h}^{\tau},a_{h}^{\tau}\in\mathcal{Z}^{k}_{h,\bar{\eta}^{\prime}}}\chi^{\tau}_{h}(f^{(n)})\Big{|}.$
@@
$\|f^{(n)}_{\bar{\eta}}-f^{(n)}_{\bar{\eta}^{\prime}}\|_{\infty}$ $=\max_{s,a}\Big{|}\sum_{n^{\prime}=1}^{n}{H^{n^{\prime}}}[\psi_{n^{\prime}}([\mathbb{P}\bar{\eta}](s,a))-\psi_{n^{\prime}}([\mathbb{P}\bar{\eta}^{\prime}](s,a))]r^{n-n^{\prime}}{/H^{n-1}}\Big{|}$
@@
$=\sum_{n=1}^{N}2(f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau}))(f^{(n)}_{\bar{\eta}}(s_{h}^{\tau},a_{h}^{\tau})-{\psi_{n}\Big{(}(\mathcal{B}_{r_{h}^{\tau}})_{\#}\mathbb{E}_{s_{h+1}^{\tau}\sim\mathbb{P}_{h}(\cdot|s_{h}^{\tau},a_{h}^{\tau})}[{\bar{\eta}}(s_{h+1}^{\tau})]\Big{)})}$
@@
$\leq{\max_{n\in[N]}\Big{\{}2(f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-\psi_{n}\Big{(}(\mathcal{B}_{r_{h}^{\tau}})_{\#}{\bar{\eta}}(s_{h+1}^{\tau})\Big{)})\Big{\}}}\sum_{n=1}^{N}\Big{|}f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})\Big{|}$
${\leq 2(H+1)\sum_{n=1}^{N}\Big{|}f^{(n)}(s_{h}^{\tau},a_{h}^{\tau})-f_{\bar{\eta}}^{(n)}(s_{h}^{\tau},a_{h}^{\tau})\Big{|}}$
@@
Lemma D.1.
Let $Q^{k}_{h}(s,a)\coloneqq{\min\{H,\tilde{f}^{k}_{h}(s,a)+b^{k}_{h}(s,a)\}}$ for some bonus function $b^{k}_{h}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. If $b^{k}_{h}(s,a)\geq w^{(1)}((\mathcal{F}^{N})^{k}_{h},s,a)$ , then with probability at least $21-\delta/$ 2,
@@
Lemma D.2 (Regret decomposition).
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\bm{1}\{{b^{k}_{h}(s^{k}_{h},a^{k}_{h})}>\epsilon\}\leq\left(\frac{4\beta(\mathcal{F}^{N},\delta)}{\epsilon^{2}}+1\right)\text{dim}_{E}(\mathcal{F}^{N},\epsilon)$
@@
$4\beta(\mathcal{F}^{N},\delta)\geq\|f-g\|^{2}_{\mathcal{Z}^{k}}{\geq\|f^{(1)}-g^{(1)}\|^{2}_{\mathcal{Z}^{k}}}\geq L\epsilon^{2}$
@@
Lemma D.4 (Refined version of Lemma 10 in Wang et al. (2020)).
@@
We first sort the sequence $\{b^{k}_{h}(s^{k}_{h},a^{k}_{h})\}_{(k,h)\in[K]\times[H]}$ in a decreasing order and denote it by $\{e_{1},\ldots,e_{T}\}(e_{1}\geq e_{2}\geq\dots\geq e_{T})$. By Lemma D.3, for any constant $M>0$ and $e_{t}\geq 1/\sqrt{M}T$, we have
@@
Conditioning on the above event and Lemma D.4, we have